OpenStudy (anonymous):

2log[2](2)+2log[2](6)-log[2](3x)=3 x=2 x=6 x=16 x=18

OpenStudy (anonymous):

2log[2](2)+2log[2](6)-log[2](3x)=3 There are a number of ways to solve this... first I see that log[2](2) = 1 2*1 +2log[2](6)-log[2](3x)=3 2log[2](6)-log[2](3x)=1 Next use nlog(a) = log(a^n) log[2](36)-log[2](3x)=1 Then log(a) - log(b) = log(a/b) log[2](36/3x) = 1 log[2](12/x) = 1 Now apply the definition of log base 2: 2^1 = 12/x 2x = 12 x=6 Should substitute into the original to verify that it works.

OpenStudy (anonymous):

thanks